Which lottery presents the player with the best odds for
winning, (A or B)? Defend your answer.

A = C(39,5)
B = C(40,4)

Answer by htmentor(1343)

The number of winning combinations for lottery A = 39!/(5!*34!)
The number of winning combinations for lottery B = 40!/(4!*36!)
Taking the ratio B/A gives:
B/A = 40!/(4!*36!)*(5!*34!)/39!
Cancelling terms and simplifying gives:
B/A = 10/63
Therefore, there are 6.3 times more winning combinations for lottery A, so it presents the best odds.